OpenStudy (anonymous):

x intercept (4,0) y intercept (0, -3) slope and equation that defines f

OpenStudy (anonymous):

m=-3-0/0-4 m=-3/-4 m=3/4 y=mx+b using (4,0) 0=(3/4)4+b 0=3+b b=-3 therefore y=(3/4)x-3